# VLM-Social-Nav: Socially Aware Robot Navigation through Scoring using Vision-Language Models

### v2

@@ IV Experiments
In this section, we detail the implementation of VLM-Social-Nav and describe our qualitative and quantitative results including comparisons with other methods and the user study.
@@ IV-A Implementation Details
VLM-Social-Nav is tested on a Turtlebot 2 equipped with a Velodyne VLP16 LiDAR, and a Zed 2i camera, and a laptop with intel i7 CPU and an Nvidia GeForce RTX 2080 GPU. We use YOLO [YOLO2023] as our real-world perception model to detect key objects (e.g., humans, door, and gesture) and Generative Pre-trained Transformer 4 with Vision (GPT-4V) [gpt4v2023] as our VLM to comprehend the social dynamics and output the immediate preferred robot action. We combined our method with a low-level motion planner DWA [fox1997dynamic]. We compare VLM-Social-Nav with DWA without social cost $\mathcal{C}_{\textrm{social}}$ and BC [pomerleau1988alvinn] method trained on a state-of-the-art, large-scale social navigation dataset, SCAND [karnan2022scand]. The dataset contains various examples of socially compliant navigation behaviors teleoperated by humans including sticking to the right of the road, waiting for a human to pass, and following a crowd.
To validate VLM-Social-Nav, we carefully follow the social robot navigation studies [francis2023principles, pirk2022protocol] that have set up the benchmark scenarios and the metrics for measuring social compliance. We present qualitative, quantitative, and user study results in four different indoor social navigation scenarios:
@@ IV-B Qualitative Result
  • Frontal Approach: The robot is expected to yield or slow down and modify its original trajectory not to obstruct the human path. Similar to driving rules in North America, it is conventional to keep on the right.
@@
Fig. 4 shows snapshots of the resulting robot motion using VLM-Social-Nav in four different scenarios. We demonstrate that VLM-Social-Nav follows the social convention and navigates toward its goal as expected. Fig. 1 illustrates the resulting trajectories of VLM-Social-Nav in comparison to those of DWA and BC methods. A notable observation is that while DWA also effectively avoids collisions with individuals, VLM-Social-Nav generates trajectories that align more closely with social norms. For instance, in the frontal approach scenario, while DWA tends to maneuver around the person either to the right or left, VLM-Social-Nav predominantly bypasses them on the right side. Similarly, in the intersection scenario, whereas DWA occasionally obstructs the person’s path by veering to avoid collision directly in front, VLM-Social-Nav endeavors to pass behind the individual. Additionally, BC successfully avoids humans but fails to recover and follow the original path. This leads to many failures in reaching the goal. The accompanying supplementary video shows the resulting robot motions.
@@ 8 The robot followed my command
9 I felt the robot paid attention to what I was doing.
@@ 10 The robot let me cross the intersection by maintaining a safe and
12 The robot paid attention to what I was doing.
@@ IV-C Quantitative Result
To further validate VLM-Social-Nav, we evaluate the methods using three different metrics [francis2023principles]. The success rate describes whether the robot reaches the goal. For the frontal approach with gesture scenario, we mark it as successful when the robot reacts to the gesture. The collision rate describes whether the robot collided with the human or other objects in the environment. We also mark it as collision when we manually intervene to avoid an imminent collision with the human subject or surroundings. The user study score is an average score we obtained from the user study detailed in Section IV-D.
Table I reports the results averaged over 21 runs for each method and scenario. The results demonstrate that VLM-Social-Nav, DWA with social cost, outperforms other methods in every metric. DWA excels at following a path smoothly, yet it faces challenges in collision avoidance as it relies solely on the LiDAR sensor and it does not consider socially compliance. Most of the collisions occurred when DWA navigated in a way that interfered with a person’s path, for example, going in front of the person when intersecting. We also observe that the outcomes of BC varied. At times, when attempting to avoid collisions, it failed to return to its original path and failed to reach the goal. Conversely, there were instances where it didn’t attempt collision avoidance at all, resulting in collisions with the participants. For gesture recognition, only our proposed method successfully responded to the participants’ gestures. In total, VLM-Social-Nav achieves at least $27.38\%$ improvement in the average success rate and $19.05\%$ improvement in the average collision rate in the four social navigation scenarios.
@@ IV-D User Study
To validate the social compliance of VLM-Social-Nav, we conduct a user study. We ask the participants to walk along the predefined trajectory and ask them to answer questionnaires about the robot motion [pirk2022protocol] (Table II). * denotes negatively formulated questions, for which we reverse-code the ratings to ensure comparability with the positively formulated ones. The three methods are randomly shuffled and repeated three times. Each scenario is tested on seven participants. We use a five-level Likert scale to ask participants to rate their agreement toward these statements.
Fig. 5 and the user study scores on Table I show the study result. The plot shows the per-question average scores for the three methods in each scenario. Based on the results, it’s evident that VLM-Social-Nav receives the highest level of agreement from participants across all questions, indicating its strong adherence to social norms. The standard error of the BC method was large, indicating that the performance of the BC method was not consistent. The score difference between VLM-Social-Nav and DWA was not large in the narrow doorway scenario. This is because, when attempting to enter the narrow doorway, the robot motion results from DWA indicated that it failed to find a plan and froze, resembling almost a complete stop as the result of VLM-Social-Nav.
Figure 5: User Study Average Scores: the per-question average scores for the three methods in each scenario. The results indicate that VLM-Social-Nav (Ours) garnered the highest level of agreement from participants across all questions, highlighting its robust alignment with social norms.
@@ IV-E Discussion
Real-time navigation with VLM: GPT-4 and similar large VLMs require several seconds to respond to prompts, making continuous querying impractical for real-time navigation tasks. To address this, we optimized VLM-Social-Nav in two ways: first, by formatting prompts and providing predefined choices, which resulted in reduced response times. Second, we minimize queries by using a perception model to detect social cues, allowing for timely VLM queries only when necessary. These choices enable and allow average response times of 2-3 seconds, sufficient for human interaction and navigation. While such a limitation can be problematic in more dynamic scenarios that require frequent interactions, future advancements in fast large language models promise further extensions of our approach.
@@
Socially aware navigation with VLM: We observe that VLMs can analyze and reason about social interactions from single images. Using various single images, including those collected by ourselves and from social robot navigation datasets [nguyenmusohu, karnan2022scand], VLMs accurately describe scenes and suggest socially compliant navigation strategies with reasons. For instance, with the image shown in Fig. 3, GPT-4V describes the scene as a person is walking towards the camera along a sidewalk. To navigate this situation, GPT-4V advises to yield the right of way because it is generally customary to keep to the right side of the path when encountering someone coming from the opposite direction, similar to driving rules. However, despite their powerful capabilities, VLMs can still make mistakes. Therefore, it is not yet safe to rely solely on VLMs for navigation decisions. Instead, we incorporate their output as a cost term in our overall decision-making process.
A single action output of VLM: Our approach begins with the observation that the robot action can be effectively translated into linguistic terms. VLMs still face challenges in accurate spatial understanding. Mapping spatial information from images to a textual format for communicating with VLMs is another challenging research problem, unless we use a fine-tuned foundation model. Therefore, adopting an approach where robot actions can be easily translated into textual output offers a straightforward yet robust solution that effectively mitigates VLM hallucination. However, this may have the drawback of not being able to plan long-range trajectories ahead. This is our immediate interest for future work; to extend our approach to planning longer-range trajectories for outdoor navigation.
@@
More challenging scenarios: Although our robot experiments were conducted only indoors, according to the example in Fig. 3, VLM-Social-Nav can be extended to outdoor scenarios in more complex environments. VLMs successfully retrieve environment information that is significant for outdoor social robot navigation, such as sidewalks, zebra crossings, and cars. This will be our immediate focus for future work, i.e., to advance our work into global outdoor navigation. We also aim to extend our approach to complex scenarios involving multiple individuals. When tested with the scenario with a group of people, VLM-Social-Nav successfully outputs socially compatible actions, realizing the group of people in the scene. However, when multiple groups are present, simple directions like left or right may not suffice to describe effective robot navigation. We plan to explore multiple trajectory generation methods [liang2023mtg, liang2024dtg] and selection techniques using visual prompts [nasiriany2024pivot, liu2024moka] to enhance interaction precision.
@@ V Conclusion
We propose a novel social navigation approach based on VLMs, focusing on real-time, socially compliant decision-making in human-centric environments. We utilize the perception model to detect important social entities and prompt a VLM to generate guidance for socially compliant behavior. VLM-Social-Nav features a VLM-based scoring that ensures socially appropriate and effective robot actions. This minimizes the dependence on extensive training datasets and eliminates the necessity for explicit rules or rewards typically associated with reinforcement learning methods. By furnishing textual instructions to VLM, we can instruct the robot to adhere to specific navigation rules, such as navigating on the right or left according to cultural norms. In practice, it results in improved socially compliant navigation in human-shared environments. We demonstrate and evaluate our system in four different real-world social navigation scenarios with a Turtlebot robot.

### v3

@@ IV Experiments
In this section, we describe the details of the implementation and qualitative and quantitative experimental results.
@@ IV-A Implementation Details
VLM-Social-Nav is tested on a Turtlebot 2 equipped with a Velodyne VLP16 LiDAR, a Zed 2i camera, and a laptop with an Intel i7 CPU and an Nvidia GeForce RTX 2080 GPU. We use YOLO [YOLO2023] as our real-world perception model to detect key objects. In our experiment, we focus on key social cues, i.e., humans, doors, and gestures, which are critical considerations when navigating in socially rich environments [kruse2013human]. Generative Pre-trained Transformer 4 with Vision (GPT-4V) [gpt4v2023] is used as our VLM to comprehend the social dynamics and output the immediate preferred robot action. This follows a preliminary study that compared its performance with other large and small VLMs. GPT-4V was able to produce reliable results with high consistency, achieving a reasonable average inference time of around three seconds. We combined our approach with a low-level motion planner DWA [fox1997dynamic]. We compare VLM-Social-Nav with DWA without social cost $\mathcal{C}_{\textrm{social}}$ and BC [pomerleau1988alvinn] trained on a state-of-the-art, large-scale social navigation dataset, SCAND [karnan2022scand]. The dataset contains various examples of socially compliant navigation behaviors teleoperated by humans including sticking to the right of the road and waiting for a human to pass. We expect that the model can learn to output socially compliant navigation behaviors like the human demonstrations.
Evaluating the social aspects of social robot navigation is inherently challenging [tsoi2024robot]. To validate VLM-Social-Nav, we carefully follow the social robot navigation studies [francis2023principles, pirk2022protocol], which set up the benchmark scenarios and the metrics for measuring social compliance. We present qualitative, quantitative, and user study results in four different social navigation scenarios:
@@
TABLE I: Quantitative Results: performance comparisons using BC [pomerleau1988alvinn], DWA [fox1997dynamic], and VLM-Social-Nav
Metric Method Scenario
Success Rate (%) $\uparrow$ BC 38.10 0 33.33 42.86
DWA 100 0 90.48 100
VLM-Social-Nav 100 100 100 100
Collision Rate (%) $\downarrow$ BC 42.86 66.67 28.57 38.10
DWA 28.57 19.05 19.05 38.10
VLM-Social-Nav 14.29 0 4.76 9.52
User Study Score $\uparrow$ BC 2.80 $\pm$ 1.45 2.23 $\pm$ 1.54 2.80 $\pm$ 1.40 2.60 $\pm$ 1.33
DWA 3.99 $\pm$ 0.80 3.38 $\pm$ 0.64 3.57 $\pm$ 0.62 3.59 $\pm$ 0.83
VLM-Social-Nav 4.31 $\pm$ 0.72 4.28 $\pm$ 0.56 4.35 $\pm$ 0.70 4.04 $\pm$ 0.74
@@ IV-B Qualitative Result
  • Frontal Approach: The robot is expected to yield or slow down and modify its original trajectory so that it does not obstruct the human path. Like driving rules in North America, it is conventional to keep on the right.
@@
Fig. 4 shows snapshots of the resulting robot motion using VLM-Social-Nav in four different scenarios. We demonstrate that VLM-Social-Nav follows the social convention and navigates toward its goal as expected. Fig. 1 illustrates the resulting trajectories of VLM-Social-Nav in comparison to those of DWA and BC methods. A notable observation is that, while DWA also effectively avoids collisions with individuals, VLM-Social-Nav generates trajectories that align more closely with social norms. For instance, in the frontal approach scenario, while DWA tends to maneuver around the person either to the right or left, VLM-Social-Nav predominantly bypasses the person on the right side. Similarly, in the intersection scenario, whereas DWA occasionally obstructs the person’s path by veering to avoid collision directly in front, VLM-Social-Nav adjusts its trajectory to pass behind the individual, adapting effectively to the human’s movement direction. Additionally, BC avoids humans but fails to recover and follow the original path. This leads to many failures in reaching the goal. The accompanying supplementary video shows the resulting robot motions.
@@ 8 The robot followed my command
9 I felt like the robot paid attention to what I was doing.
@@ 10 The robot let me cross the intersection by maintaining a safe and
12 I felt like the robot paid attention to what I was doing.
@@ IV-C Quantitative Result
To further validate VLM-Social-Nav, we evaluate the methods using three different metrics. The success rate describes whether the robot reaches the goal. For the frontal approach with gesture scenario, we mark it as successful when the robot reacts to the gesture. The collision rate describes whether the robot collided with the human or other objects in the environment. We also mark it as in collision when we manually intervene to avoid an imminent collision with the human subject or surroundings. The user study score is an average score we obtained from the user study detailed in Section IV-D.
Table I reports the results averaged over 21 runs for each method and scenario. The results demonstrate that VLM-Social-Nav, DWA with social cost, outperforms other methods in every metric. DWA excels at following a path smoothly, yet it faces challenges in collision avoidance as it relies solely on the LiDAR sensor and does not consider social compliance. Most of the collisions occurred when DWA navigated in a way that interfered with a person’s path, for example, going in front of the person when intersecting. We also observe that the outcomes of BC varied. At times, when attempting to avoid collisions, it failed to return to its original path and failed to reach the goal. Conversely, there were instances where it didn’t attempt collision avoidance at all, resulting in collisions with the participants. For gesture recognition, only our proposed method successfully responded to the participants’ gestures. VLM-Social-Nav improves the average success rate by $27.38\%$ and reduces the average collision rate by $19.05\%$ across four social navigation scenarios.
@@ IV-D User Study
To validate the social compliance of VLM-Social-Nav, we conduct a user study. We ask the participants to walk along the predefined trajectory and then to answer questionnaires about the robot motion [pirk2022protocol] (Table II). * denotes negatively formulated questions, for which we reverse-code the ratings to ensure comparability with the positively formulated ones. The three methods are randomly shuffled and repeated three times. Each scenario is tested on seven participants. We use a five-level Likert scale to ask participants to rate their agreement with these statements.
Fig. 5 and the user study scores in Table I show the study result. The plot shows the per-question average scores for the three methods in each scenario. Based on the results, it’s evident that VLM-Social-Nav receives the highest level of agreement from participants across all questions, indicating its strong adherence to social norms. The standard error of the BC method was large, indicating that the performance of the BC method was not consistent. The score difference between VLM-Social-Nav and DWA was not large in the narrow doorway scenario. This is because, when attempting to enter the narrow doorway, DWA often failed to find a plan and froze, resembling the result of VLM-Social-Nav, a complete stop in front of the doorway.
Figure 5: User Study Average Scores: the per-question average scores for the three methods in each scenario. The results indicate that VLM-Social-Nav earned the highest level of agreement from participants across all questions, highlighting its robust alignment with social norms.
@@ IV-E Discussion
Real-time navigation with VLM: GPT-4V and similar large VLMs require several seconds to respond to prompts, making continuous querying impractical for real-time navigation tasks. To address this, we optimized VLM-Social-Nav in two ways: first, by formatting prompts and providing predefined choices, which resulted in reduced response times. Second, we minimize queries by using a perception model to detect social cues, allowing for timely VLM queries only when necessary. These choices enable and allow average response times of 2-3 seconds, sufficient for human interaction and navigation. While such a limitation can be problematic in more dynamic scenarios that require frequent interactions, future advancements in fast large language models promise further extensions of our approach.
@@
Socially aware navigation with VLM: We observe that VLMs can analyze and reason about social interactions from single images. Using various single images, including those collected by ourselves and from social robot navigation datasets [nguyenmusohu, karnan2022scand], VLMs accurately describe scenes and suggest socially compliant navigation strategies with reasons. For instance, for the image shown in Fig. 3, GPT-4V describes the scene as a person is walking towards the camera along a sidewalk. To navigate this situation, GPT-4V advises the robot to yield the right of way because it is generally customary to keep to the right side of the path when encountering someone coming from the opposite direction, similar to driving rules. However, despite their powerful capabilities, VLMs can still make mistakes. Therefore, relying solely on VLMs for navigation decisions is not safe. Instead, we incorporate their output as a cost term in our overall decision-making process.
@@
More challenging scenarios: Although our robot experiments were conducted only indoors, according to the example in Fig. 3, VLM-Social-Nav can be extended to outdoor scenarios in more complex environments. VLMs successfully retrieve significant environmental information for outdoor social robot navigation, such as sidewalks, zebra crossings, and cars. This will be our immediate focus for future work, i.e., to advance into global outdoor navigation. We also aim to extend our approach to complex scenarios involving multiple individuals. When tested with the scenario with one person, VLM-Social-Nav successfully outputs socially compatible actions, realizing the group of people in the scene. However, when multiple groups are present, simple directions like left or right may not suffice to describe effective robot navigation.
@@ V Conclusion
We propose a novel social navigation approach based on VLMs, focusing on real-time, socially compliant decision-making in human-centric environments. We utilize the perception model to detect important social entities and prompt a VLM to generate guidance for socially compliant behavior. VLM-Social-Nav features a VLM-based scoring that ensures socially appropriate and effective robot actions. This minimizes the dependence on extensive training datasets and eliminates the necessity for explicit rules or hand-tuned parameters typically associated with imitation learning approaches. By furnishing textual instructions to VLM, we can instruct the robot to adhere to specific navigation rules, such as navigating on the right or left according to cultural norms. However, interpreting social rules and deriving appropriate actions from them based on raw robot perception remains complex. VLMs interpret social situations and determine actions based on these rules, offering a more nuanced approach than simpler rule- or planning-based methods. We demonstrate and evaluate our system in four different real-world social navigation scenarios with a Turtlebot robot.
One immediate future work is to explore visual prompting methods [nasiriany2024pivot, liu2024moka] to enhance spatial reasoning in VLMs by marking the images. Another promising future direction is to explore open-source VLMs, such as LLaVA [liu2024visual]. Their access to lower-level information, such as log probabilities could help detect and address hallucinations. It is also interesting to develop a VLM that generates high-level social navigation instructions through chain-of-thought reasoning, and integrate it into an autonomous navigation system.